OpenStudy (anonymous):

Find the derivative of the function. y = [(x2 + 4)/(x2 − 4)]^3

OpenStudy (anonymous):

is it x^2 or 2x

OpenStudy (anonymous):

\[(\frac{x ^{2}+4 }{x ^{2}-4 })^{3}\]

OpenStudy (anonymous):

there are three ways to do this question which one did you try

OpenStudy (anonymous):

quotient rule

OpenStudy (anonymous):

\[3(\frac{x ^{2}+4 }{x ^{2}-4 })^{2}\]

zepdrix (zepdrix):

So that is the derivative of the outer function. Looks good so far. Now you multiply by the derivative of the inside? So what'd you get for your quotient rule? :D

OpenStudy (anonymous):

\[3(\frac{x ^{2}+4 }{x ^{2}-4 })^{2}\]*\[\frac{ (2x)(x ^{2}-4)-(x ^{2}+4)(2x) }{ (x ^{2}-4)^{2} }\]

OpenStudy (anonymous):

is it correct?

OpenStudy (cwrw238):

yes - now you need to multiply by the earlier result

OpenStudy (anonymous):

do i have to expand?

OpenStudy (cwrw238):

no but check if you can simplify